OpenStudy (anonymous):

The ratio between two quantities is 2 : 7. If the second quantity is 9.9 kg, find the first. A) 34.65 kg B) 2.83 kg C) 2.2 kg D) 2.92 kg

OpenStudy (anonymous):

Hi again :D

OpenStudy (anonymous):

you there

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So whats this one!

OpenStudy (anonymous):

Question is too vague.

OpenStudy (anonymous):

Mean to comfusing

OpenStudy (anonymous):

I think it's A.

OpenStudy (anonymous):

Ratio of first quantity to second quantity is constant 2/7 = y/9.9, where y represents the first quantity The cross products of proportions

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

This is what i can help on

OpenStudy (anonymous):

It's b then.

OpenStudy (anonymous):

when i do it it gives me 2.82857

OpenStudy (anonymous):

and alot of wacky numberrs

OpenStudy (anonymous):

Ya, It would be B

OpenStudy (nubeer):

Ratio 2 : 7 now we make the second side 1 as it's data is given so divide both sides by 7 2/7 : 1 now for second quantity 9.9 is given so multiply both sides with 9.9 9.9 * (2/7) : 1 * 9.9

OpenStudy (anonymous):

For the set of ratios, find the two that are proportional. 49 , 7, 45 182 26 160 A) 49/182 and 7/26 B) 7/26 and 45/160 C) 49/182 and 45/160

OpenStudy (anonymous):

How about this one

OpenStudy (anonymous):

I'll let nubeer take over from here.

OpenStudy (anonymous):

NUUU

OpenStudy (anonymous):

I like you zordoloom O:

OpenStudy (anonymous):

I never got a best response.

OpenStudy (anonymous):

i give you one :D

OpenStudy (anonymous):

See ;D

OpenStudy (nubeer):

lol @zordoloom it's all fine you can explain i was just helping thats all :)

OpenStudy (anonymous):

Thanks nubeer

OpenStudy (anonymous):

Well nbeer can you help me

OpenStudy (anonymous):

For the set of ratios, find the two that are proportional. 49 , 7, 45 182 26 160 A) 49/182 and 7/26 B) 7/26 and 45/160 C) 49/182 and 45/160

OpenStudy (nubeer):

its A option 49/182 = 7/26

OpenStudy (anonymous):

Write a proportion that could be used to solve for the variable s. Then solve. 5 pounds of seed for 2000 square feet x pounds of seed for 1500 square feet A) 5/1500 = x/2000; 7.5 B) 5/1500 = x/2000; 8.5 C) 5/2000 = x/1500; 8.5 D) 5/2000 = x/1500; 3.75

OpenStudy (anonymous):

And this one

OpenStudy (anonymous):

You there

OpenStudy (nubeer):

it's C for this one. 5/2000 = x/1500 x = 3.75

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

i know this might be to much asking can you help me still?

OpenStudy (nubeer):

yeah sure.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Solve the proportion. y - 4 = y + 2 3 5 A) 12 B) 11 C) 13 D) 52

OpenStudy (anonymous):

Inbewtween

OpenStudy (anonymous):

the the nubmers

OpenStudy (anonymous):

are a fraction

OpenStudy (anonymous):

i got an idea

OpenStudy (nubeer):

hmm idea? and what is that/

OpenStudy (anonymous):

There

OpenStudy (anonymous):

In the attachment is the question

OpenStudy (nubeer):

5(y-4) =3(y+2) open brackets and multiply 5y-20 = 3y +6 now just find the value of y.. can you?

OpenStudy (anonymous):

Ummm

OpenStudy (anonymous):

is it 3.25

OpenStudy (nubeer):

nop.. can you put steps so i can tell where u made mistake.,

OpenStudy (anonymous):

i jsut did the other you gave me

OpenStudy (anonymous):

jsut added the like terms

OpenStudy (anonymous):

added the 20 on both sides

OpenStudy (anonymous):

divided it

OpenStudy (anonymous):

gave me 3.5

OpenStudy (nubeer):

divided by?

OpenStudy (anonymous):

26

OpenStudy (nubeer):

noo.. let me explain 5y-20=3y+6 first add 20 on both sides.

OpenStudy (anonymous):

26

OpenStudy (nubeer):

5y = 3y +26

OpenStudy (anonymous):

ok

OpenStudy (nubeer):

now subtract -3y on both sides.

OpenStudy (anonymous):

2y

OpenStudy (nubeer):

yes so 2y = 26 now divide both sides by 2

OpenStudy (anonymous):

oh 13

OpenStudy (anonymous):

zordoloom is on :D

OpenStudy (nubeer):

yes.. :)

OpenStudy (anonymous):

Julia surveyed 45 students about their preferred lunch. Thirty preferred chicken. Half of those students preferred chicken with barbecue sauce. What percent should Julia report as preferring chicken with barbecue sauce? A) 33.66% B) 32.34% C) 33.33% D) 31.35%

OpenStudy (anonymous):

Help on dis!

OpenStudy (nubeer):

ok out of 45, 30 like chicken.

OpenStudy (anonymous):

huh?

OpenStudy (nubeer):

and half of 30 like with sauce so 30/2 15 students like with sauce.

OpenStudy (anonymous):

so what percent

OpenStudy (nubeer):

so (15/45)*100%

OpenStudy (anonymous):

33.3333333333333333

OpenStudy (anonymous):

Thats the answer

OpenStudy (nubeer):

yes.

OpenStudy (anonymous):

Its not on options O:

OpenStudy (anonymous):

Nvm found it

OpenStudy (anonymous):

The actual length of a dust mite is 0.3 millimeters. The model length is 3 cm. What is the scale of the model? A) 100:1 B) 1: 100 C) 10:1 D) 1:10

OpenStudy (nubeer):

hmm yes C one.

OpenStudy (anonymous):

And this one

OpenStudy (nubeer):

C for this one too.

OpenStudy (anonymous):

so 10:1

OpenStudy (nubeer):

yes

OpenStudy (anonymous):

The number of students at Franklin High School increased from 840 to 910 over a 5-year period. What was the percent of change? A) 9.22% B) 7.29% C) 8.33% D) 6.55%

OpenStudy (anonymous):

And this one?

OpenStudy (nubeer):

(910-840)/840 * 100%

OpenStudy (anonymous):

15.476 and bunch of whacky numbers i got

OpenStudy (nubeer):

doubt that.. u did it wrong.. try again (910-840)/840 = ?

OpenStudy (anonymous):

0.083333333333333

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Thats what i got

OpenStudy (nubeer):

now this answeer multiply 100

OpenStudy (nubeer):

so it's 8.33

OpenStudy (anonymous):

Oo

OpenStudy (anonymous):

Yay still zordolloom is here :D